OpenStudy (anonymous):

Python 3.4.1 Help Please: I will post my code that i have gotten so far below, What i need to do is if given x amount of inches, and a cost (per yard), how do i find the total cost? I know i have to convert it to yards but just having a brain fart. Please help me out here

OpenStudy (anonymous):

``` total = eval(input('How many inches of material? ')) cost = eval(input('What is the cost (per yard)? ')) age = eval(input('What is your age? ')) #Start Math yards = total // 36 total = total % 36 feet = total // 12 total = total % 12 inches = total #End Math print('You bought ' + str(yards) + ' yards, ' + str(feet) + ' feet ' + str(inches) + ' inches of material') #Start Math Part II #End Math Part II print('Material Cost: ') ```

OpenStudy (anonymous):

im going into my next class now, even if im not online please just post what your willing to to help me, ill be back soon.

OpenStudy (espex):

Take your inches, we are assuming linear, and divide it by 36 to get the inches to yard conversion. \[\frac{total \space inches}{1} \times \frac{1\space yard}{36\space inches} = \frac{yards}{1}\] Then multiply your cost per yard for the total. \[\frac{total \space yard}{1} \times \frac{dollars}{1 \space yard} = \frac{dollars}{1}\]